I figured out (a) already so I'm looking for help on (b). Also, in which direction will the positively charged protein feel a force due to the DNA backbone? A vector needs to be drawn.
Question &: The backbone of DNA is highly negatively charged. This net charge arises from the phosphate groups along the backbone of each chain, where each phosphate group has a net charge of -le (a) Consider the phosphate groups to be separated by 3.4 A. What is the force between adjacent groups? Is it attractive or repulsive? (b) Now, imagine a protein is nearby, which has a +le charge that is 5 A away from the DNA backbone, and is situated directly between two phosphate groups as shown. Below is a sketch of the geometry of the problem, labeling the charges and the distances between them. Determine the angle θ.
media%2F52e%2F52ec3490-b36a-4a7a-8144-14
Answer #1

5A

from the figure we represented 4 points A,B,C,D

tan\theta in a right angled triangle is given by

  tan\theta = side opposite to \theta / side adjacent to \theta

=> tan\theta=\frac{BC}{AB}

but   BC=\frac{DC}{2}

=>   tan\theta =\frac{DC}{2(AB)}

=> tan\theta =\frac{3.4}{2(5)}

=> tan\theta =0.34

=>   \theta=tan^{-1}(0.34)

=> \theta=18.78^{o}